OpenStudy (anonymous):

Find the potential energy of a 40.0 kg ball lodged in a tree 10.0 m high

OpenStudy (anonymous):

3920 Joule...

OpenStudy (anonymous):

E=mgh

OpenStudy (anonymous):

Thank you .

OpenStudy (anonymous):

neva mind..

OpenStudy (anonymous):

Would the answer be 400 J ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you forgot to multiply g

OpenStudy (anonymous):

1/2 ?

OpenStudy (anonymous):

do you know what gravitational acceleration is?

OpenStudy (anonymous):

Kind of yes and no .

OpenStudy (anonymous):

Nevermind. I got it.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

@isadrawr..the gravitation potential energy has an expression sumthing like this:- E=mgh where m is mass ...h is hieght....and g is accelaration due to gravity... g=9.8m/s near the sea level.. thus E=40*9.8*10=3920 joules...

OpenStudy (anonymous):

Thank you .

OpenStudy (anonymous):

E=mgh , g=9.8\[\rightarrow\]E=40*9.8*10=3920j